OpenStudy (nastech):

find (dz/dr) and (dr/ds) if z=sin^-1(y/x)+2^(xln(y)) and x=(r^2+s^2)^(1/2) and y=log base r of s^2

OpenStudy (anonymous):

what is LXIX in roman numerals HELP !!!!! I REALLY NEED HELP and PLEASE EXPLAiN HOW U GET THE ANSWER

OpenStudy (zehanz):

@keefe10: Did you accidently type in your question as an anwer to Nastech's?

OpenStudy (anonymous):

Sorry, @Hoa I was just keeping you entertained :) You might want to use this notation instead \[\huge \frac{\partial z}{\partial x}\cdot\frac{\partial x}{\partial r}\]\[\huge \frac{\partial z}{\partial y}\cdot\frac{\partial y}{\partial r}\] It's \partial :)

OpenStudy (anonymous):

@PeterPan Continue have fun with what you think it's fun. Fooling others is not my habit. Good luck

OpenStudy (anonymous):

There really is no need to take things so personally (and so seriously) In fact, all I asked you to do was define and contrast partial derivatives and stuff, which is rather harmless. If you're still mad at me, then I really am sorry :) If you can't forgive me, then that's out of my hands :>

OpenStudy (anonymous):

what kind of saying sorry is yours? ok, let it go. I am not mad at you anymore.

OpenStudy (anonymous):

:)